OpenStudy (goformit100):

Write whether the square of any positive integer can be of the form 3m+ 2, where m is a natural number. Justify your answer.

OpenStudy (goformit100):

@genius12 Sir Please HELP me

mathslover (mathslover):

Goformit, do you know euclid's division algorithm?

OpenStudy (goformit100):

ya , but how to use it here ?

mathslover (mathslover):

It says, any positive integer a can be re-written as : \(\bf{a=bq+r}\) (a is divided by b, having quotient q and remainder r ) . - Short definition for euclid's division algorithm.

OpenStudy (goformit100):

ok

OpenStudy (anonymous):

If m is a natural number it belongs to the set N = {0, 1 , 2, 3...} Let's try prove this statement through contradiction, i.e proof by contradiction. To build this proof, we must contradict the statement, i.e. we must assume that 3m + 2 is not the square of any positive integer where m is a natural number.\[\bf 3m+2 \ne x^2, \ x \in \mathbb{Z}^+ \]Choosing a value for m, such as 5 gives us:\[\bf 17 \ne x^2, \ x \in \mathbb{Z}^+\]And this is true since 17 is not a perfect square. @goformit100

mathslover (mathslover):

Now, say the number is a. When, a is divided by 3, I can write it as: \(\mathsf{a=3q+r}\) Note that, remainder can not exceed 3 and will not be negative. \(\mathsf{0 \le r < 3}\) So, possible cases are : \(\mathsf{a =3q + 0\\ a = 3q + 1\\ a = 3q + 2}\)

OpenStudy (anonymous):

So since the contradiction is correct, the original statement must be false. Hence the square of any positive integer cannot be of the form 3m+ 2, where m is a natural number.

mathslover (mathslover):

Square it now: Case 1 : a = 3q \(a^2 = 9q^2\) \(a^2 = 3(3q^2) \) \(q^2\) is also an integer. Let us say that, \(3q^2\) = m So, \(a^2 = 3m\)

mathslover (mathslover):

@goformit100 , is your question correctly stated here? Re-check it.

mathslover (mathslover):

Case : 2 \(\mathsf{ a = 3q + 1\\ a^2 = (3q+1)^2 \\ a^2 = 9q^2 + 1 + 6q \\ a^2 = 9q^2 + 6q + 1\\ a^2 = 3(3q^2 + 2) + 1 \\ a^2 = 3m + 1}\) Case : 3 \(\mathsf{a = 3q + 2\\ a^2 = (3q+2)^2\\ a^2 = 9q^2 + 4 + 12q \\ a^2 = 3(3q^2 + 4q + 1) + 1\\ a^2 = 3m+1}\) Square of any positive integer can be in the form of 3m or 3m+1. But not 3m+2.

OpenStudy (zarkon):

a little quicker would be to ... \[a=3q+m\] \[a^2=9q^2+6mq+m^2\] \[a^2=3(3q^2+2mq)+m^2\] and then notice that \(m^2\cancel{\equiv}2\text{ mod }3\) \[m=0,1,2\]

mathslover (mathslover):

@Zarkon , Nice method. Thanks for introducing it to us.

mathslover (mathslover):

@goformit100 , If the questions asks, that whether the given statement holds true for any positive integer, then NO! It is not correct.

mathslover (mathslover):

But , @Zarkon , unfortunately, I am not good at congruent modulo. Can you teach me some of the basics?

OpenStudy (zarkon):

i'll just write out what we have above \[0^2\equiv 0 (mod 3)\] \[1^2\equiv 1 (mod 3)\] \[2^2\equiv 1 (mod 3)\]

OpenStudy (goformit100):

Thanks

mathslover (mathslover):

Okay. I think, I will have to study it from any youtube lecture. Thanks @Zarkon ...

OpenStudy (zarkon):

you could have also done cases on my last statement...what if m=0 or 1 or 2...sorta like you did

OpenStudy (zarkon):

have you taken discrete math?

OpenStudy (zarkon):

it is usually introduced in that class

mathslover (mathslover):

Oh! I am in high school and we have to study all type of mathematics we have (not all but those who are of our level) ... So, no division, only study and study all topics.

mathslover (mathslover):

And Olympiad is also near, so I have to completely understand congruent modulo for solving tricky problems.

OpenStudy (zarkon):

ic...good luck with that

mathslover (mathslover):

Thanks :)